Grade 12th passAlgebra

Simplify and solve the following linear equations.
0.25(4f – 3) = 0.05(10f – 9)

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Harshit Singh
5 Years ago
Dear Student

0.25(4f–3) = 0.05(10f–9)
⇒f-0.75=0.5f-0.45
⇒f-0.5f=-0.45+0.75
⇒ 0.5f = 0.30
⇒f = 0.30/0.5
⇒f=3/5
⇒f=0.6

Thanks